# THOMAS et al

### v1

@@ 4.2 Data Simulation
The statistical models that included random effects at the individual level ("GP.m" and "brms.m") treated the third covariate $x_{3}$ as representing a measurement at the individual-level data, for the sake of comparison, but this choice is not expected to have a large influence on the results, as there was no variation in the simulated data beyond the observed covariates and shared noise.
@@ 4.3 Losses and Testing
The models were evaluated using 10-fold cross-validation to predict the test outputs given the training data and test covariates. Loss functions were evaluated using the mean predictive $f_{pm}$ across posterior samples and the observed data, using an absolute/L1 loss for Gaussian outputs $y_{g}$ and for binomial outputs $y_{b}$, the logarithm of one minus the probability of the observed data ("the log probability of the wrong answer"), i.e. $l(y_{g},f_{pm})=|y_{g}-f_{pm}|$ and $l(y_{b},f_{pm})=\log(\Phi(-f_{pm}*(2y_{b}-1))$, where $\Phi$ is the probit link.
@@ 4.4 Results for Simulated Data
We see the predictive loss results in Table 1 and Figure 3. We see that the brms model effectively fails and returns samples from the prior on binomial output 1 and Gaussian output 1, which are derived from the same latent parametric function. The GP models consequently registered significantly lower losses with large effect sizes. We see a small and possibly spurious effect of the GP models appearing to not predict around the prior at $p=0.5$. For the second continuous output, we again see the GP models outperforming the brms models, with the GP.f performing substantially better than GP.m. For the second binomial output, we see that the brms models get many of the labels correct with high confidence, but also many of the labels wrong with high confidence. Consequently, the results here are more mixed, with the only significant differences being brms.m outperforming both GP.m and brms.f. We would expect from the absolute value present in 11 that the smooth GP model would find the second and fourth endpoints more challenging to model accurately.
@@
1 GP.m - -1.000 -1.000 -0.102 - -0.778 -0.778 0.077
brms.f <.001 - 0.062 1.000 <.001 - -0.005 0.769
brms.m <.001 0.439 - 1.000 <.001 0.953 - 0.769
GP.f 0.405 <.001 $0$ 0.664 <.001 <.001 -
2 GP.m - -0.567 -0.567 0.588 - -0.083 0.251 -0.074
brms.f <.001 - 0.084 0.790 0.892 - 0.598 0.018
brms.m <.001 0.291 - 0.794 0.008 <.001 - -0.196
GP.f <.001 <.001 <.001 - 0.892 0.892 0.055 -
@@ 4.5 Observational Time-Series Data
We see the predictive loss results for the first real data set in Table 2 and Figure 4. We see that all of the methods were confidently correct in their predictions of labels for the binomial outputs, suggesting that this is a relatively easy classification problem. That said, when considering the ranking of the predictions, we see that GP.m model outperforms the other methods on the binomial outputs. We see a small number of extremely confident predictions from the brms model, especially for output 3, which is a concern even if they are correct.
When considering the continuous output, we see more mixed results between the methods, with the only clearly significant result being that the GP.f model appears to perform uniformly worse than the other methods.
@@
G.1 GP.m - 0.057 0.064 -0.359 - -0.530 -0.435 -0.839 GP.m B.2
brms.f 0.356 - 0.134 -0.405 <.001 - 0.840 -0.061 brms.f
brms.m 0.356 0.015 - -0.408 <.001 <.001 - -0.442 brms.m
GP.f <.001 <.001 <.001 - <.001 0.200 <.001 - GP.f
B.1 GP.m - -0.941 -0.687 -0.750 - -0.482 -0.176 -0.879 GP.m B.3
brms.f <.001 - 0.955 0.390 $0$ 0.987 -0.010 brms.f
brms.m <.001 $0$0.098 0.000 <.001 - -0.686 brms.m
GP.f <.001 <.001 0.038 - <.001 0.835 <.001 - GP.f
@@ 4.6 Randomised Control Trial Data
1 GP.m - -0.598 -0.599 -0.447 - -0.974 -0.776 -0.197
brms.f <.001 - -0.014 0.134 <.001 - 0.989 0.859
brms.m <.001 0.423 - 0.135 <.001 <.001 - 0.503
GP.f <.001 <.001 <.001 - <.001 <.001 <.001 -
2 GP.m - -0.358 -0.360 -0.356 - -0.940 -0.680 0.417
brms.f <.001 - -0.088 -0.008 <.001 - 0.881 0.817
brms.m <.001 <.001 - -0.005 <.001 <.001 - 0.643
GP.f <.001 1.000 1.000 - <.001 <.001 <.001 -
@@ 5 Discussion
In this work, we have demonstrated that the repeated measurements used in applied medical research can be used to scale complex Bayesian nonparametric Gaussian Process models to practical clinical questions. The more complex models have showed to have increased predictive ability compared to the widely-used parametric models, indicating that this may be worthwhile, if achieving optimal model specification is a concern.
@@
Acknowledgments
Author contributions
Financial disclosure
None reported.
Conflict of interest
The authors declare no potential conflict of interests.
Appendix A Stan Outputs
@@
HU.Intercept -1.39 0.074 -1.537 -1.247 1 2409.083 2984.331
Peter.Intercept -23.136 15.093 -62.433 -6.119 1.003 608.858 576.982
Sumit.Intercept -23.363 18.786 -77.043 -5.53 1.006 409.13 320.44
Thien.Intercept -29.083 19.913 -70.655 -7.379 1.003 691.947 498.643
HU.Sex -0.254 0.051 -0.357 -0.154 1.001 4559.867 2846.231
HU.Age -0.082 0.022 -0.125 -0.039 1.001 4361.788 2682.417
HU.KbyV 0.15 0.028 0.094 0.207 1.001 3720.045 3164.562
HU.TubeVoltage -0.007 0.061 -0.125 0.114 1 3883.938 3248.508
HU.Side -0.034 0.043 -0.12 0.05 1 4765.954 2693.787
HU.Times.1 0.772 0.08 0.615 0.926 1.001 2144.028 2807.039
HU.Times.2 1.699 0.082 1.537 1.866 1.001 2219.165 2908.287
HU.Times.3 2.093 0.083 1.932 2.262 1 2269.678 2760.094
HU.Times.4 2.161 0.08 2.003 2.318 1.001 2205.402 2842.623
HU.Times.5 2.16 0.081 1.998 2.32 1.001 2183.258 2429.128
HU.Times.6 2.159 0.081 2.003 2.317 1 2267.362 3046.82
Peter.Sex -1.596 0.392 -2.383 -0.842 1.001 3919.9 3341.905
Peter.Age -0.625 0.151 -0.931 -0.34 1.002 5417.444 2985.744
Peter.KbyV 0.296 0.19 -0.07 0.679 1 3325.893 2899.162
Peter.TubeVoltage 1.505 0.421 0.662 2.33 1.002 2808.877 2631.265
Peter.Side 0.101 0.283 -0.439 0.659 1 5410.548 2982.479
Peter.Times.1 19.842 15.096 2.695 58.589 1.003 609.021 557.101
Peter.Times.2 23.304 15.094 6.205 62.54 1.003 608.571 545.979
Peter.Times.3 25.541 15.108 8.451 64.9 1.003 607.598 567.206
Peter.Times.4 25.885 15.102 8.738 65.16 1.002 613.898 549.396
Peter.Times.5 26.028 15.118 8.906 65.348 1.003 611.03 581.704
Peter.Times.6 26.01 15.096 8.902 65.252 1.003 612.908 577.387
Sumit.Sex -1.364 0.325 -2.022 -0.746 1 4423.775 2895.507
Sumit.Age -0.551 0.127 -0.793 -0.303 1.001 5916.021 2742.276
Sumit.KbyV 0.5 0.154 0.205 0.808 1.001 3755.939 2738.673
Sumit.TubeVoltage 1.049 0.349 0.38 1.754 1.001 3319.896 2957.351
Sumit.Side -0.034 0.231 -0.495 0.415 1 5586.167 2965.208
Sumit.Times.1 20.573 18.783 2.707 74.057 1.006 408.385 317.588
Sumit.Times.2 23.32 18.79 5.438 76.846 1.006 408.887 320.442
Sumit.Times.3 25.238 18.792 7.402 79.115 1.006 409.265 319.583
Sumit.Times.4 25.295 18.788 7.376 78.9 1.005 409.093 321.446
Sumit.Times.5 25.529 18.783 7.659 78.986 1.005 409.948 320.44
Sumit.Times.6 24.87 18.781 7.055 78.473 1.006 409.119 321.843
Thien.Sex -1.093 0.516 -2.116 -0.103 1.001 3390.629 2920.511
Thien.Age -0.562 0.208 -0.988 -0.157 1 5525.843 2864.83
Thien.KbyV -0.035 0.276 -0.577 0.509 1 3796.754 3311.912
Thien.TubeVoltage 2.026 0.607 0.854 3.222 1 3172.331 2965.172
Thien.Side 0.195 0.405 -0.585 0.973 1.001 5236.692 2325.445
Thien.Times.1 24.506 19.934 2.887 66.063 1.003 691.319 479.517
Thien.Times.2 28.927 19.911 7.254 70.355 1.003 692.155 495.169
Thien.Times.3 32.396 19.948 10.587 74.065 1.003 689.624 489.518
Thien.Times.4 34.038 19.995 11.864 75.923 1.003 691.37 485.938
Thien.Times.5 34.031 19.989 11.909 76.187 1.003 695.89 504.097
Thien.Times.6 33.999 19.993 11.874 76.105 1.003 683.468 502.573
@@
HU.Intercept -1.387 0.105 -1.591 -1.18 1.001 1618.397 2350.325
Peter.Intercept -33.07 17.474 -76.833 -11.798 1.001 610.753 403.092
Sumit.Intercept -29.19 28.557 -$1$8.163 1.033 96.551 22.218
Thien.Intercept -47.712 25.873 -119.681 -16.076 1.005 397.244 235.653
HU.Sex -0.249 0.104 -0.457 -0.045 1 1274.419 1887.114
HU.Age -0.083 0.045 -0.171 0.006 1.001 1298.837 1541.561
HU.KbyV 0.152 0.055 0.042 0.258 1.003 1179.928 2072.619
HU.TubeVoltage -0.003 0.12 -0.239 0.235 1.001 1153.473 1620.713
HU.Side -0.03 0.038 -0.104 0.044 1.007 8636.042 2727.371
HU.Times.1 0.772 0.069 0.636 0.906 1.001 3222.682 3355.595
HU.Times.2 1.701 0.071 1.563 1.837 1 3227.142 3142.511
HU.Times.3 2.092 0.072 1.95 2.233 1 3132.795 3264.571
HU.Times.4 2.16 0.07 2.019 2.301 1.001 3122.661 2327.567
HU.Times.5 2.16 0.07 2.024 2.303 1.001 3183.713 3003.711
HU.Times.6 2.15 0.072 2.007 2.29 1.001 3253.032 3026.83
Peter.Sex -4.558 1.801 -8.348 -1.288 1.001 1423.471 2296.226
Peter.Age -1.553 0.759 -3.166 -0.149 1 1302.35 2167.791
Peter.KbyV 0.033 0.909 -1.802 1.771 1.004 1102.953 1896.869
Peter.TubeVoltage 3.637 1.948 -0.117 7.662 1.001 1250.605 1622.67
Peter.Side 0.315 0.414 -0.483 1.116 1.001 8134.444 3079.231
Peter.Times.1 25.238 17.369 4.68 68.967 1.002 610.631 399.844
Peter.Times.2 34.946 17.552 13.457 79.709 1.001 603.05 408.878
Peter.Times.3 40.177 17.659 18.292 85.207 1.001 598.898 406.589
Peter.Times.4 40.714 17.667 18.804 85.767 1.001 600.787 407.753
Peter.Times.5 40.987 17.666 19.086 85.603 1.001 604.068 428.653
Peter.Times.6 40.957 17.673 19.018 85.762 1.001 612.928 424.599
Sumit.Sex -2.38 1.079 -4.695 -0.294 1.001 1356.179 1981.976
Sumit.Age -0.942 0.444 -1.834 -0.11 1.003 1391.182 2059.489
Sumit.KbyV 0.857 0.56 -0.212 1.956 1.003 1385.784 2115.462
Sumit.TubeVoltage 1.7 1.175 -0.576 4.029 1.002 1272.342 1796.238
Sumit.Side -0.081 0.301 -0.675 0.496 1.001 7380.028 2808.069
Sumit.Times.1 24.668 28.559 3.78 137.087 1.031 104.278 22.229
Sumit.Times.2 29.354 28.572 8.362 141.328 1.031 103.682 22.201
Sumit.Times.3 32.853 28.572 11.76 144.49 1.032 103.688 22.119
Sumit.Times.4 32.896 28.569 11.864 144.481 1.031 103.616 22.181
Sumit.Times.5 33.322 28.568 12.247 144.342 1.032 103.926 22.156
Sumit.Times.6 32.2 28.571 11.09 144.183 1.032 103.205 22.143
Thien.Sex -2.909 2.095 -7.371 0.794 1.001 1832.914 2101.764
Thien.Age -1.337 0.879 -3.253 0.247 1.001 1666.775 2237.486
Thien.KbyV -0.32 1.08 -2.498 1.775 1.004 1420.407 2315.322
Thien.TubeVoltage 5.034 2.561 0.484 10.854 1 1559.872 1907.257
Thien.Side 0.637 0.608 -0.522 1.839 1.003 6508.089 2826.776
Thien.Times.1 36.157 25.34 5.81 108.251 1.004 402.642 236.309
Thien.Times.2 48.473 25.996 16.357 121.175 1.004 391.273 239.821
Thien.Times.3 55.615 26.395 22.403 128.826 1.004 389.505 234.506
Thien.Times.4 57.812 26.517 24.452 130.267 1.005 390.179 235.505
Thien.Times.5 57.797 26.51 24.464 130.758 1.004 393.287 238.189
Thien.Times.6 57.725 26.516 24.307 131.551 1.004 391.34 230.294
@@
Appendix B Random effects in the Intrinsic Coregionalization Model

### v2

@@ 4.2 Data Simulation
The statistical models that included random effects at the individual level (”GP.m” and ”brms.m”) treated the third covariate $x_{3}$ as representing a measurement at the individual-level data, for the sake of comparison, but this choice is not expected to have a large influence on the results, as there was no variation in the simulated data beyond the observed covariates and shared noise.
@@ 4.3 Losses and Testing
The models were evaluated using 10-fold cross-validation to predict the test outputs given the training data and test covariates. Loss functions were evaluated using the mean predictive $f_{pm}$ across posterior samples and the observed data, using an quadratic/L2 loss for Gaussian outputs $y_{g}$, while for binomial outputs $y_{b}$, we used the logarithm of one minus the probability of the observed data (”the log probability of the wrong answer”), i.e. $l(y_{g},f_{pm})=(y_{g}-f_{pm})^{2}$ and $l(y_{b},f_{pm})=\log(\Phi(-f_{pm}*(2y_{b}-1))$, where $\Phi$ is the Gaussian cumulative distribution function.
@@ 4.4 Results for Simulated Data
We see the predictive loss results in Table 1 and Figure 3. We see that the brms model effectively fails and returns samples from the prior on binomial output 1 and Gaussian output 1, which are derived from the same latent parametric function. The GP models consequently registered significantly lower losses with large effect sizes. We see a small and possibly spurious effect of the GP models appearing to not predict around the prior at $p=0.5$. For the second continuous output, we again see the GP models outperforming the brms models, with the GP.f performing substantially better than GP.m. For the second binomial output, we see that the brms models get many of the labels correct with high confidence, but also many of the labels wrong with high confidence. Consequently, the results here are more mixed, with the only significant differences being brms.m outperforming both GP.m and brms.f. We would expect from the absolute value present in 9 that the smooth GP model would find the second and fourth endpoints more challenging to model accurately, relative to the first and third.
@@
1 GP.m - -1.000 -1.000 -0.091 - -0.778 -0.778 0.077
brms.f <.$<.001$ 0.044 1.000 <.$<.001$0.015 0.769
brms.m <.$<.001$ 0.582 - 1.000 <.$<.001$ 0.852 - 0.769
GP.f 0.508 <.$<.001$ <.$<.001$ 0.664 <.$<.001$ <.$<.001$
2 GP.m - -0.608 -0.602 0.534 - -0.083 0.255 -0.074
brms.f <.$<.001$ 0.182 0.777 0.892 - 0.606 0.018
brms.m <.$<.001$ 0.022 - 0.781 0.007 <.$<.001$0.197
GP.f <.$<.001$ <.$<.001$ <.$<.001$ 0.892 0.892 0.053 -
@@ 4.5 Observational Time-Series Data
We see the predictive loss results for the first real data set in Table 2 and Figure 4. We see that all of the methods were confidently correct in their predictions of labels for the binomial outputs, suggesting that this is a relatively easy classification problem. That said, when considering the ranking of the predictions, we see that GP.m model outperforms the other methods on the binomial outputs, the GP.f model mostly outperforms the brms models, and the brm.m model outperforms the brms.f model. We see a small number of extremely confident predictions from the GP.m model, especially for output 1, which is a concern even if they are correct, as the probabilities are unlikely to be well-calibrated.
When considering the continuous output, we see more mixed results between the methods, with the only clearly significant result being that the GP.m model performs uniformly better than the other methods, whose losses are otherwise indistinguishable.
@@
G.1 GP.m - -0.390 -0.389 -0.343 - -0.878 -0.859 -0.839 GP.m B.2
brms.f <.$<.001$ 0.014 -0.005 <.$<.001$ 0.885 0.552 brms.f
brms.m <.$<.001$ 1.000 - -0.006 <.$<.001$ <.$<.001$ 0.327 brms.m
GP.f <.$<.001$ 1.000 1.000 - <.$<.001$ <.$<.001$ <.$<.001$ GP.f
B.1 GP.m - -0.954 -0.953 -0.750 - -0.965 -0.837 -0.879 GP.m B.3
brms.f <.$<.001$ 0.962 0.833 <.$<.001$ 0.987 0.312 brms.f
brms.m <.$<.001$ <.$<.001$ 0.737 <.$<.001$ <.$<.001$0.106 brms.m
GP.f <.$<.001$ <.$<.001$ <.$<.001$ <.$<.001$ <.$<.001$ <.$<.001$ GP.f
@@ 4.6 Randomised Control Trial Data
1 GP.m - -0.593 -0.593 -0.444 - -0.974 -0.774 -0.197
brms.f <.$<.001$0.017 0.107 <.$<.001$ 0.989 0.859
brms.m <.$<.001$ 0.352 - 0.108 <.$<.001$ <.$<.001$ 0.501
GP.f <.$<.001$ <.$<.001$ <.$<.001$ <.$<.001$ <.$<.001$ <.$<.001$
2 GP.m - -0.357 -0.359 -0.360 - -0.940 -0.681 0.417
brms.f <.$<.001$0.058 -0.007 <.$<.001$ 0.878 0.817
brms.m <.$<.001$ 0.003 - -0.004 <.$<.001$ <.$<.001$ 0.644
GP.f <.$<.001$ 1.000 1.000 - <.$<.001$ <.$<.001$ <.$<.001$
@@ 5 Discussion
In this work, we have demonstrated that the repeated measurements used in applied medical research can be used to scale complex Bayesian nonparametric Gaussian Process models to practical clinical questions. The more complex nonparametric models have showed to have increased predictive ability compared to the widely-used parametric models, with one of the two GP models considered demonstrating the best predictive ability in every output of the real data sets considered. Consequently, pursuing the GP modelling framework is advantageous, if achieving the best possible model specification is a concern.
@@
{appendices}
6 Stan Outputs
@@
HU.Intercept -1.219 0.081 -1.376 -1.062 1.001 7872.147 2868.837
Peter.Intercept -4.675 0.483 -5.666 -3.759 1 6087.325 2978.142
Sumit.Intercept -3.235 0.354 -3.939 -2.575 1 7279.6 3308.639
Thien.Intercept -9.156 0.981 -11.182 -7.371 1.001 3760.612 3083.354
HU.Sex -0.262 0.063 -0.388 -0.137 1 6183.659 3057.013
HU.Age -0.085 0.027 -0.137 -0.031 1 8696.971 2677.263
HU.KbyV 0.161 0.035 0.093 0.229 1.001 5213.191 3076.073
HU.TubeVoltage 0 0.075 -0.146 0.15 1 4978.226 3690.191
HU.Side -0.033 0.053 -0.133 0.07 1.001 7692.8 2804.257
HU.Time 2.455 0.094 2.269 2.638 1.002 9716.016 2633.583
Peter.Sex -1.144 0.311 -1.759 -0.531 1 4677.549 3330.772
Peter.Age -0.481 0.134 -0.748 -0.219 1 7339.025 2992.232
Peter.KbyV 0.252 0.165 -0.08 0.575 1 5266.446 3241.451
Peter.TubeVoltage 1.184 0.364 0.48 1.906 1.001 4342.362 3319.535
Peter.Side 0.082 0.25 -0.42 0.575 1 8018.375 2380.159
Peter.Time 9.8 0.774 8.344 11.35 1 5192.483 3105.538
Sumit.Sex -1.078 0.261 -1.586 -0.582 1 5927.317 3224.886
Sumit.Age -0.432 0.111 -0.649 -0.217 1 8482.177 3230.506
Sumit.KbyV 0.45 0.138 0.177 0.714 1.001 5359.511 3409.304
Sumit.TubeVoltage 0.851 0.306 0.258 1.445 1 4939.906 3211.057
Sumit.Side -0.022 0.213 -0.429 0.388 1.001 8566.185 2964.559
Sumit.Time 6.276 0.485 5.367 7.223 1.001 7182.14 3646.269
Thien.Sex -0.817 0.451 -1.704 0.059 1 5344.679 3392.033
Thien.Age -0.464 0.187 -0.831 -0.107 1 7095.642 3507.823
Thien.KbyV -0.02 0.238 -0.489 0.443 1.002 5797.254 3370.95
Thien.TubeVoltage 1.695 0.542 0.693 2.787 1 4477.558 3326.391
Thien.Side 0.156 0.361 -0.544 0.882 1 6439.037 2641.699
Thien.Time 20.262 1.988 16.651 24.36 1 3534.26 2875.124
@@
HU.Intercept -1.229 0.11 -1.44 -1.005 1.001 1883.872 2635.57
Peter.Intercept -6.888 0.929 -8.859 -5.181 1 2155.894 2448.99
Sumit.Intercept -4.592 0.74 -6.093 -3.181 1.003 2059.998 2829.259
Thien.Intercept -12.612 1.799 -16.488 -9.523 1.003 1913.781 2247.858
HU.Sex -0.261 0.107 -0.471 -0.053 1 1326.375 2245.068
HU.Age -0.083 0.045 -0.173 0.004 1.003 1412.836 2148.39
HU.KbyV 0.159 0.055 0.049 0.264 1.003 1510.37 2118.139
HU.TubeVoltage 0.006 0.121 -0.233 0.238 1.002 1436.848 2183.957
HU.Side -0.03 0.049 -0.125 0.066 1 5286.117 3292.53
HU.Time 2.45 0.086 2.281 2.617 1.001 5186.381 2701.354
Peter.Sex -1.657 0.802 -3.306 -0.14 1.001 1390.627 2088.921
Peter.Age -0.688 0.332 -1.355 -0.03 1.003 1265.427 1992.546
Peter.KbyV 0.303 0.395 -0.459 1.077 1.002 1455.29 2067.752
Peter.TubeVoltage 1.715 0.882 -0.029 3.455 1.003 1361.107 2431.833
Peter.Side 0.146 0.304 -0.452 0.747 1.003 5842.318 2887.942
Peter.Time 14.326 1.286 11.963 17.022 1 2992.62 2513.319
Sumit.Sex -1.457 0.71 -2.848 -0.063 1 1349.48 2298.628
Sumit.Age -0.586 0.291 -1.164 -0.026 1.002 1313.955 2027.691
Sumit.KbyV 0.585 0.353 -0.133 1.274 1.003 1336.028 1903.154
Sumit.TubeVoltage 1.132 0.788 -0.466 2.695 1.001 1316.225 1989.443
Sumit.Side -0.021 0.253 -0.521 0.481 1.001 5200.054 2922.916
Sumit.Time 8.754 0.713 7.429 10.248 1.003 3797.535 3110.568
Thien.Sex -1.15 0.857 -2.879 0.547 1 1805.37 2631.778
Thien.Age -0.64 0.366 -1.367 0.049 1.001 1739.675 2030.872
Thien.KbyV -0.1 0.44 -0.981 0.751 1.002 1957.174 2417.188
Thien.TubeVoltage 2.423 1.008 0.498 4.525 1.002 1914.17 2449.821
Thien.Side 0.281 0.42 -0.521 1.119 1.001 5454.051 2826.627
Thien.Time 27.753 3.479 21.682 35.474 1.003 1881.15 2370.816
@@
7 Random effects in the Intrinsic Coregionalization Model
@@
8 Competing interests
The authors declare no potential conflict of interests.
9 Author contributions statement
10 Acknowledgments